# Pilot Contamination-Aware Graph Attention Network for Power Control in CFmMIMO

### v1

@@ Abstract
Optimization-based power control algorithms are predominantly iterative with high computational complexity, making them impractical for real-time applications in cell-free massive multiple-input multiple-output (CFmMIMO) systems. Learning-based methods have emerged as a promising alternative, and among them, graph neural networks (GNNs) have demonstrated their excellent performance in solving power control problems. However, all existing GNN-based approaches assume ideal orthogonality among pilot sequences for user equipments (UEs), which is unrealistic given that the number of UEs exceeds the available orthogonal pilot sequences in CFmMIMO schemes. Moreover, most learning-based methods assume a fixed number of UEs, whereas the number of active UEs varies over time in practice. Additionally, supervised training necessitates costly computational resources for computing the target power control solutions for a large volume of training samples. To address these issues, we propose a graph attention network for downlink power control in CFmMIMO systems that operates in a self-supervised manner while effectively handling pilot contamination and adapting to a dynamic number of UEs. Experimental results show its effectiveness, even in comparison to the optimal accelerated projected gradient method as a baseline.
Index Terms:
Graph attention network, power control, pilot contamination, cell-free massive MIMO.
@@ I Introduction
To fully realize the benefits of CFmMIMO, several challenges must be addressed. In particular, power control [^1] We focus on downlink power control in this paper. and pilot allocation are critical. With an increasing number of UEs, especially in CFmMIMO systems, the limited number of orthogonal pilot sequences becomes insufficient, leading to pilot contamination. Pilot contamination significantly degrades the communication quality, and thus, its impact cannot be overlooked when addressing power control problems. Traditional power control algorithms are predominantly optimization-based, with computational complexity increasing sharply as the numbers of APs and UEs grow. For CFmMIMO systems with a large number of APs and UEs, existing algorithms lead to high computational costs and processing time, which runs counter to the demands for high efficiency and low latency.
To address this challenge, learning-based algorithms have been proposed to compute power allocation coefficients with significantly improved efficiency. Various neural network models, including convolutional neural networks [4, 5], transformer neural networks [6, 7], and graph neural networks (GNNs) [8, 9, 10, 11, 12] have been explored using supervised or self-supervised training approaches. GNNs have demonstrated excellent performance in solving power control problems thanks to their structural similarity to wireless communication networks [9, 10]. As mentioned earlier, pilot contamination cannot be ignored when addressing power control. However, to the best of our knowledge, all existing GNN-based methods assume the use of ideally orthogonal pilot sequences, which is a big stretch especially in CFmMIMO systems. To bridge this gap, we propose a pilot contamination-aware graph attention network to handle power control in CFmMIMO systems. The proposed network model utilizes a GNN enhanced with attention mechanism [6, 7] to achieve power control with the awareness of pilot contamination, and is able to handle dynamic scenarios with a varying number of UEs without requiring retraining. Besides, the network is trained in a self-supervised manner, reducing the workload of computing target power control coefficients. Experimental results are presented to compare the proposed method against the first-order accelerated projected gradient (APG) method described in [13] and transformer in [7].
@@ II-A Cell-Free Massive MIMO
Consider the downlink power control in a CFmMIMO system, where all UEs are served by APs using common time-frequency resources, and all APs are synchronized. The system operates in time-division duplex mode. The system consists of $M$ APs and $K$ single-antenna UEs randomly distributed over a large coverage area, with each AP equipped with $N$ antennas. Assuming that channels are block-fading and remain constant over a coherence block of $T_{\rm c}$ symbols, the uplink channel from the $k$th UE to the $m$th AP can be expressed as
$\mathbf{g}_{mk}=\sqrt{\beta_{mk}}{\mathbf{h}_{mk}},$ (1)
where $\beta_{mk}$ denotes the large-scale fading coefficient, and $\mathbf{h}_{mk}\in\mathbb{C}^{N}$ denotes the small-scale coefficient vector for an $N$-antenna array. The entries of $\mathbf{h}_{mk}$ are independent and identically distributed (i.i.d.) complex Gaussian random variables with zero mean and unit variance, i.e., $\mathcal{CN}(0,1)$.
[…]
@@
$\mathbf{Y}_{{\rm up},m}=\sqrt{\zeta_{\rm p}T_{\rm p}}\sum_{i=1}^{K}\mathbf{g}_{mi}\bm{\psi}_{i}^{\rm H}+\mathbf{W}_{{\rm up},m},$ (2)
where $\zeta_{\rm p}$ denotes the normalized transmit signal-to-noise ratio per pilot symbol, $\left(\cdot\right)^{\rm H}$ represents the conjugate transpose operation, and $\mathbf{W}_{{\rm up},m}\in\mathbb{C}^{N\times T_{\rm p}}$ is the noise matrix whose entries follow i.i.d. $\mathcal{CN}(0,1)$. The minimum mean square error (MMSE) estimate of the channel between the $m$th AP and the $k$th UE can be derived as [13]
$\hat{\mathbf{g}}_{mk}=\frac{\sqrt{\zeta_{\rm p}T_{\rm p}}\beta_{mk}}{1+\zeta_{\rm p}T_{\rm p}\sum_{i=1}^{K}\beta_{mi}\left|\bm{\psi}_{i}^{\rm H}\bm{\psi}_{k}\right|^{2}}\mathbf{Y}_{{\rm up},m}\bm{\psi}_{k},$ (3)
[…]
$\overline{g}_{mk}=\mathbb{E}\left[\left|\hat{\mathbf{g}}_{mk}[n]\right|^{2}\right]=\frac{\zeta_{\rm p}T_{\rm p}\beta_{mk}^{2}}{1+\zeta_{\rm p}T_{\rm p}\sum_{i=1}^{K}\beta_{mi}\left|\bm{\psi}_{i}^{\rm H}\bm{\psi}_{k}\right|^{2}}.$ (4)
APs cooperate to deal with power control with the knowledge of large-scale fading information through the central processing unit, while each AP works independently to apply a beamforming scheme with the estimated small-scale fading coefficients. Denote the power control coefficient matrix as $\mathbf{M}\in\mathbb{R}^{M\times K}$, and the $(m,k)$-th matrix element, $\mu_{mk}$, is the power control coefficient for the signal transmitted from the $m$th AP to the $k$th UE. According to the practical constraint on the total transmit power per AP, the feasible set of power control coefficients can be expressed as
$\mathcal{S}=\left\{\mathbf{M}\mid\mu_{mk}\geq 0;\left\|\bm{\mu}_{m}\right\|^{2}\leq\frac{1}{N},\forall m,k\right\},$ (5)
@@
$\mathbf{x}_{m}=\sqrt{\zeta_{\rm d}}\sum_{k=1}^{K}\frac{\mu_{mk}}{\sqrt{\overline{g}_{mk}}}\hat{\mathbf{g}}_{mk}^{*}s_{k},$ (6)
@@
$\gamma_{k}=\frac{\zeta_{\rm d}\left(\overline{\bm{\mu}}_{k}^{\rm T}\bm{\nu}_{kk}\right)^{2}}{\sum_{\begin{subarray}{c}i=1\\ i\neq k\end{subarray}}^{K}\zeta_{\rm d}\left(\overline{\bm{\mu}}_{i}^{\rm T}\bm{\nu}_{ik}\right)^{2}+\frac{\zeta_{\rm d}}{N}\sum_{i=1}^{K}\left\|\overline{\mathbf{B}}_{k}\overline{\bm{\mu}}_{i}\right\|^{2}+\frac{1}{N^{2}}},$ (7)
@@
$\bm{\nu}_{ik}[m]=\left|\bm{\psi}_{i}^{\rm H}\bm{\psi}_{k}\right|\sqrt{\overline{g}_{mi}}\frac{\beta_{mk}}{\beta_{mi}}.$ (8)
@@ II-B Optimization Problem Formulation
Define $\mathbf{B}\in\mathbb{R}_{+}^{M\times K}$ as the large-scale fading coefficient matrix with $\beta_{mk}$ as the $(m,k)$-th element, and let $\bm{\Phi}\in\mathbb{R}^{K\times K}$ denote the pilot allocation matrix, where the $(i,j)$-th element is given by $\left|\bm{\psi}_{i}^{\rm H}\bm{\psi}_{j}\right|$. Here $\mathbb{R}_{+}^{M\times K}$ represents the set of matrices of dimension $M\times K$ with non-negative real entries. From (7) and (9), SE depends on both the power control coefficients $\mathbf{M}$ and pilot allocation matrix $\bm{\Phi}$, given a large-scale fading matrix $\mathbf{B}$. When $T_{\rm p}\geq K$, pilot sequences are orthogonal across all UEs, making their impact on SE negligible. However, in massive MIMO scenarios, $T_{\rm p}<K$ often occurs. Therefore, incorporating the information of pilot assignment is essential for optimizing power control under pilot contamination.
@@
By adopting a smoothing technique with smoothing parameter $\lambda$, the objective function in (10) can be reformulated as [13]
@@
Define a deep neural network as a map, $\mathbf{M}=f(\bm{\Phi},\mathbf{B};\mathbf{\Theta})$. Here $f$ maps $\bm{\Phi}$ and $\mathbf{B}$ to the power control matrix $\mathbf{M}$, and $\mathbf{\Theta}$ represents the trainable weights of the network. Our goal is to determine the optimal weights $\mathbf{\Theta}_{\rm opt}$ such that the neural network generates the power control coefficient matrix $\mathbf{M}$ to maximize the objective function $u$ in (11).
@@ III-A Graph Representation
Consider a heterogeneous graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ composed of a set of nodes $\mathcal{V}$ and edges $\mathcal{E}$. Each node $v$ represents an AP-UE pair and is mapped by $\varphi(m,k)=v$, where $\varphi$ is the mapping function from an AP-UE pair to its corresponding node, and $v=M(k-1)+m$. To enable a single power control network capable of handling a varying number of UEs, $K$ is set as the maximum number of UEs, $K_{\rm max}$, within the coverage area. If the actual number of UEs, $K_{\rm act}$, is smaller than $K_{\rm max}$, an extra $K_{\rm max}-K_{\rm act}$ UEs are added as padding. To differentiate between actual and padded UEs, each node is assigned a binary attribute, ${\rm Mask}$, and ${\rm Mask}_{i}=1$ indicates that the UE associated with the $i$th node is served, while ${\rm Mask}_{i}=0$ denotes a padded UE. The bidirectional edge between the $i$th and $j$th nodes is denoted as $e_{ij}$ for $i\neq j$. Edges are categorized based on whether the connected nodes share the same AP or UE, or neither:
  • AP-type edges ($e_{ij}^{\rm AP}$): Nodes share the same AP.
  • UE-type edges ($e_{ij}^{\rm UE}$): Nodes share the same UE.
@@
To incorporate pilot contamination awareness, the pilot allocation matrix $\bm{\Phi}$ is included as an attribute for AP-type edges, capturing the orthogonality of allocated pilot sequences among UEs that share the same APs. Specifically, the $(i,j)$-th element of $\bm{\Phi}$, denoted as $\psi_{ij}$, is assigned to those AP-type edges whose two connected nodes correspond to the $i$th and $j$th UEs. Suppose there are $T$ attention layers in the graph structural network, and the input and output features of the $i$th node at the $t$th graph attention layer are represented by $\mathbf{x}_{i}(t-1)$ and $\mathbf{x}_{i}(t)$, respectively, where $t=1,\ldots,T$. The characteristics of the proposed graph structure, including the node representation $\mathbf{x}_{i}(t)$, the node attribute ${\rm Mask}_{i}$, the AP-type edges with pilot contamination information $\psi_{ij}$, and the UE-type edges, are illustrated in Fig. 1. This structure enables the network to effectively model the relationship between APs and UEs while accounting for pilot contamination.

### v4

@@ Abstract
Optimization-based power control algorithms are predominantly iterative with high computational complexity, making them impractical for real-time applications in cell-free massive multiple-input multiple-output (CFmMIMO) systems. Learning-based methods have emerged as a promising alternative, and among them, graph neural networks (GNNs) have demonstrated their excellent performance in solving power control problems. However, all existing GNN-based approaches assume ideal orthogonality among pilot sequences for user equipments (UEs), which is unrealistic given that the number of UEs exceeds the available orthogonal pilot sequences in CFmMIMO schemes. Additionally, supervised training necessitates costly computational resources for computing the target power control solutions for a large volume of training samples. To address these issues, we propose a graph attention network for downlink power control in CFmMIMO systems that operates in a self-supervised manner while effectively handling pilot contamination and adapting to a dynamic number of UEs. Experimental results show its effectiveness, even in comparison to the optimal accelerated projected gradient method as a baseline.
@@ I Introduction
To fully realize the benefits of CFmMIMO, several challenges must be addressed. In particular, power control [^1] We focus on downlink power control in a centralized fashion in this paper. and pilot allocation are critical. With an increasing number of UEs, especially in CFmMIMO systems, the limited number of orthogonal pilot sequences becomes insufficient, leading to pilot contamination. Pilot contamination significantly degrades the communication quality, and thus, its impact cannot be overlooked when addressing power control problems. Traditional power control algorithms are predominantly optimization-based, with computational complexity increasing sharply as the numbers of APs and UEs grow. For CFmMIMO systems with a large number of APs and UEs, existing algorithms lead to high computational costs and processing time, which runs counter to the demands for high efficiency and low latency.
To address this challenge, learning-based algorithms have been proposed to compute power allocation coefficients with significantly improved efficiency. Various neural network models, including convolutional neural networks [4, 5], transformer neural networks [6, 7], and graph neural networks (GNNs) [8, 9, 10, 11, 12, 13] have been explored. GNNs have demonstrated excellent performance in solving power control problems thanks to their structural similarity to wireless communication networks [9, 10]. As mentioned earlier, pilot contamination cannot be ignored when addressing power control. However, to the best of our knowledge, all existing GNN-based methods assume the use of ideally orthogonal pilot sequences, which is a big stretch, especially in CFmMIMO systems. To bridge this gap, we propose a pilot contamination-aware graph attention network to handle power control in CFmMIMO systems. The proposed network model utilizes an attention mechanism [6, 7] enhanced GNN to achieve power control with the awareness of pilot contamination in dynamic scenarios with a varying number of UEs. Besides, the network is trained in a self-supervised manner, reducing the workload of computing target power control coefficients. Experimental results are presented to compare the proposed method against the first-order accelerated projected gradient (APG) method described in [14] and transformer in [7].
@@ II-A Cell-Free Massive MIMO
Consider the downlink power control in a CFmMIMO system, where all UEs are served by APs using common time-frequency resources, and all APs are synchronized. Stringent synchronization among UEs and APs in both time and frequency domains is essential to enable coherent joint transmission in CFmMIMO architectures. Achieving such synchronization is technically challenging, especially at large scale and with distributed APs, and requires dedicated algorithms and network infrastructure. Several practical approaches have been developed to address this challenge. In cell-free architectures, synchronization is often maintained via a combination of centralized processing at the control unit, high-capacity fronthaul links, and regular synchronization signals exchanged between the APs and the central unit. Solutions can include over-the-air synchronization procedures, network-based protocols, and periodic calibration, sometimes leveraging GPS or precise network timing protocols (e.g., IEEE 1588 Precision Time Protocol). Experimental and field trials have demonstrated the feasibility of achieving tight synchronization in such distributed settings. Further advancements on this topic in terms of scalability and cost are still required [1].
The system operates in time-division duplex (TDD) mode. The system consists of $M$ APs and $K$ single-antenna UEs randomly distributed over a large coverage area, with each AP equipped with $N$ antennas. Assuming that channels are block-fading and remain constant over a coherence block of $T_{\rm c}$ symbols, the uplink channel from the $k$th UE to the $m$th AP can be expressed as
$\mathbf{g}_{m,k}=\sqrt{\beta_{m,k}}{\mathbf{h}_{m,k}},$ (1)
where $\beta_{m,k}$ denotes the large-scale fading coefficient, and $\mathbf{h}_{m,k}\in\mathbb{C}^{N}$ denotes the small-scale coefficient vector for an $N$-antenna array. The entries of $\mathbf{h}_{m,k}$ are independent and identically distributed (i.i.d.) complex Gaussian random variables with zero mean and unit variance, i.e., $\mathcal{CN}(0,1)$.
Pilot sequences consisting of $T_{\rm p}$ ($T_{\rm p}\ll T_{\rm c}$) symbols are used for channel estimation during the uplink training stage. Channel reciprocity is assumed, enabling downlink channel estimation using uplink pilot sequences. Let the pilot sequence for the $k$th UE be $\sqrt{T_{\rm p}}\bm{\varphi}_{k}\in\mathbb{C}^{T_{\rm p}}$ with $\|\bm{\varphi}_{k}\|^{2}=1$, where $\|\cdot\|$ represents the Euclidean norm of a vector. In CFmMIMO systems, where a large number of UEs are served simultaneously, the condition $T_{\rm p}<K$ often arises due to the limited length of pilot sequences, resulting in pilot contamination.
@@
$\mathbf{Y}_{m}^{\rm up}=\sqrt{\zeta_{\rm p}T_{\rm p}}\sum_{i=1}^{K}\mathbf{g}_{m,i}\bm{\varphi}_{i}^{\rm H}+\mathbf{W}_{m}^{\rm up},$ (2)
where $\zeta_{\rm p}$ denotes the normalized transmit signal-to-noise ratio per pilot symbol, $\left(\cdot\right)^{\rm H}$ represents the conjugate transpose operation, and $\mathbf{W}_{m}^{\rm up}\in\mathbb{C}^{N\times T_{\rm p}}$ is the noise matrix whose entries follow i.i.d. $\mathcal{CN}(0,1)$. The minimum mean square error (MMSE) estimate of the channel between the $m$th AP and the $k$th UE can be derived as [14]
$\hat{\mathbf{g}}_{m,k}=\frac{\sqrt{\zeta_{\rm p}T_{\rm p}}\beta_{m,k}}{1+\zeta_{\rm p}T_{\rm p}\sum_{i=1}^{K}\beta_{m,i}\left|\bm{\varphi}_{i}^{\rm H}\bm{\varphi}_{k}\right|^{2}}\mathbf{Y}_{m}^{\rm up}\bm{\varphi}_{k},$ (3)
where $\left|\cdot\right|$ denotes the absolute value. The mean square of each entry of $\hat{\mathbf{g}}_{m,k}$ is
$\overline{g}_{m,k}=\mathbb{E}\left[\left|\hat{\mathbf{g}}_{m,k}[n]\right|^{2}\right]=\frac{\zeta_{\rm p}T_{\rm p}\beta_{m,k}^{2}}{1+\zeta_{\rm p}T_{\rm p}\sum_{i=1}^{K}\beta_{m,i}\left|\bm{\varphi}_{i}^{\rm H}\bm{\varphi}_{k}\right|^{2}}.$ (4)
APs cooperate to deal with power control using large-scale fading information, which is assumed to be available through the central processing unit. Accurate large-scale fading information is critical for power allocation in CFmMIMO systems. In practice, large-scale fading coefficients vary slowly over time when UEs are stationary or moving at slow speeds, so the channel can be modeled as block-fading. These coefficients can be obtained through channel estimation using periodic pilot transmission and uplink-downlink reciprocity in TDD systems. For higher-mobility UEs, one can employ tracking algorithms (e.g., Kalman filters) along with robust estimation techniques over successive coherence intervals to track the variations of large-scale fading [16, 17], thus providing real‑time large‑scale fading estimates. The use of small-scale channel coefficients for power control is constrained by two key factors. First, small-scale fading varies rapidly over time, requiring frequent updates. Relying on it would impose significant computational overhead, making it impractical for large-scale network-level optimization tasks. Second, unlike beamforming, power control is primarily influenced by the large-scale channel characteristics; the impact of small-scale fading on power control solutions is relatively limited. Therefore, even though the small-scale channel estimates are available, only the large-scale information is used in most CFmMIMO power control [18, 19, 20].
Denote the power control coefficient matrix as $\mathbf{M}\in\mathbb{R}^{M\times K}$, and the $(m,k)$-th matrix element, $\mu_{m,k}$, is the power control coefficient for the signal transmitted from the $m$th AP to the $k$th UE. According to the practical constraint on the total transmit power per AP, the feasible set of power control coefficients can be expressed as
$\mathcal{S}=\left\{\mathbf{M}\mid\mu_{m,k}\geq 0;\left\|\bm{\mu}_{m}\right\|^{2}\leq\frac{1}{N},\forall m,k\right\},$ (5)
@@
$\mathbf{x}_{m}=\sqrt{\zeta_{\rm d}}\sum_{k=1}^{K}\frac{\mu_{m,k}}{\sqrt{\overline{g}_{m,k}}}\hat{\mathbf{g}}_{m,k}^{*}s_{k},$ (6)
@@
$\gamma_{k}=\frac{\zeta_{\rm d}\left(\overline{\bm{\mu}}_{k}^{\rm T}\bm{\nu}_{k,k}\right)^{2}}{\sum_{\begin{subarray}{c}i=1\\ i\neq k\end{subarray}}^{K}\zeta_{\rm d}\left(\overline{\bm{\mu}}_{i}^{\rm T}\bm{\nu}_{i,k}\right)^{2}+\frac{\zeta_{\rm d}}{N}\sum_{i=1}^{K}\left\|\overline{\mathbf{B}}_{k}\overline{\bm{\mu}}_{i}\right\|^{2}+\frac{1}{N^{2}}},$ (7)
@@
$\bm{\nu}_{i,k}[m]=\left|\bm{\varphi}_{i}^{\rm H}\bm{\varphi}_{k}\right|\sqrt{\overline{g}_{m,i}}\frac{\beta_{m,k}}{\beta_{m,i}}.$ (8)
@@ II-B Optimization Problem Formulation
Define $\mathbf{B}\in\mathbb{R}_{+}^{M\times K}$ as the large-scale fading coefficient matrix with $\beta_{m,k}$ as the $(m,k)$-th element, and let $\bm{\Phi}\in\mathbb{R}^{K\times K}$ denote the pilot allocation matrix, where the $(i,j)$-th element is given by $\left|\bm{\varphi}_{i}^{\rm H}\bm{\varphi}_{j}\right|$. Here $\mathbb{R}_{+}^{M\times K}$ represents the set of matrices of dimension $M\times K$ with non-negative real entries. From (7) and (9), SE depends on both the power control coefficients $\mathbf{M}$ and pilot allocation matrix $\bm{\Phi}$, given a large-scale fading matrix $\mathbf{B}$. When $T_{\rm p}\geq K$, pilot sequences are orthogonal across all UEs, making their impact on SE negligible. However, in massive MIMO scenarios, $T_{\rm p}<K$ often occurs. Therefore, incorporating the information of pilot assignment is essential for optimizing power control under pilot contamination.
@@
By adopting a smoothing technique with smoothing parameter $\lambda$, the objective function in (10) can be reformulated as [14, 15]
@@
Define a deep neural network as a map, $\mathbf{M}=f(\bm{\Phi},\mathbf{B};\mathbf{\Theta})$. Here $f$ maps $\bm{\Phi}$ and $\mathbf{B}$ to the power control matrix $\mathbf{M}$, and $\mathbf{\Theta}$ represents the trainable weights of the network. Our goal is to determine the optimal weights $\mathbf{\Theta}_{\rm opt}$ using a self-supervised learning method, in which the neural network is trained on unlabeled data, thus eliminating the computational burden of generating target power control solutions, for example, via conventional optimization algorithms. In this framework, the network learns the mapping between its input and power control coefficients $\mathbf{M}$ with the aim of maximizing the objective function in (11).
@@ III-A Graph Representation
Consider a heterogeneous graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ composed of a set of nodes $\mathcal{V}$ and edges $\mathcal{E}$. Each node $v$ represents an AP-UE pair and is mapped by $\psi(m,k)=v$, where $\psi$ is the mapping function from an AP-UE pair to its corresponding node, and v=M(k1)+mv=M(k-1)+m. The bidirectional edge between the $i$th and $j$th nodes is denoted as $e_{i,j}$ for $i\neq j$. Edges are categorized based on whether the connected nodes share the same AP or UE, or neither:
  • AP-type edges ($e_{i,j}^{\rm AP}$): Nodes share the same AP.
  • UE-type edges ($e_{i,j}^{\rm UE}$): Nodes share the same UE.
@@
To incorporate pilot contamination awareness, the pilot allocation matrix $\bm{\Phi}$ is included as an attribute for AP-type edges, capturing the orthogonality of allocated pilot sequences among UEs that share the same APs. Specifically, the $(k_{1},k_{2})$-th element of $\bm{\Phi}$, denoted by $\overline{\varphi}_{k_{1},k_{2}}$, is assigned to those AP-type edges connecting nodes corresponding to the $k_{1}$-th and $k_{2}$-th UEs. To align with the GNN framework, $\varphi_{i,j}=\overline{\varphi}_{k_{1},k_{2}}$ is introduced to represent the contamination information between nodes $i$ and $j$, which correspond to $k_{1}$-th and $k_{2}$-th UEs, respectively. Suppose there are $T$ attention layers in the graph structural network, and the input and output features of the $i$th node at the $t$th graph attention layer are represented by $\mathbf{x}_{i}(t-1)$ and $\mathbf{x}_{i}(t)$, respectively, where $t=1,\ldots,T$. The characteristics of the proposed graph structure, including the node representation $\mathbf{x}_{i}(t)$, the AP-type edges with $\varphi_{i,j}$, and the UE-type edges, are illustrated in Fig. 1. This structure enables the network to effectively model the relationship between APs and UEs while accounting for pilot contamination.